OpenStudy (anonymous):

Evaluate the expression: 5 + 3 × (5 - 2)^2

Directrix (directrix):

5 + 3 × (5 - 2)^2 = 5 + 3 x 3*3 = You do this step, okay?

OpenStudy (anonymous):

5+3*(5-2)^(2) Subtract 2 from 5 to get 3. 5+3*(3)^(2) Expand the exponent (2) to the expression. 5+3*(3^(2)) Squaring a number is the same as multiplying the number by itself (3*3). 5+3*(3*3) Multiply 3 by 3 to get 9. 5+3*(9) Multiply 3 by (9) to get 3(9). 5+3(9) Multiply 3 by the 9 inside the parentheses. 5+3*9 Multiply 3 by 9 to get 27. 5+27 Add 27 to 5 to get 32. 32

OpenStudy (anonymous):

5 + 3 × (5 - 2)^2 replace x with a * and paste to google 5 + 3 * (5 - 2)^2

OpenStudy (anonymous):

32